OpenStudy (anonymous):

find an equation for the line tangent to the curve at the given point y= (x^4 +2)/(x^2) at x=-1

OpenStudy (anonymous):

When you don't have a 0/0 you can most of the time just simply use the substitution method.

OpenStudy (anonymous):

which gives you lim= 3

OpenStudy (anonymous):

whoops wrong question lol give me a second

OpenStudy (anonymous):

Do you know how to use the quotient rule?

OpenStudy (anonymous):

Yes that is what we just learned but I am a little confused on how to apply it in this scenario

OpenStudy (anonymous):

Do you know how to find the d/dx?

OpenStudy (anonymous):

would the d/dx be 2x-(4/x^3) ?

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

so then do i need to just plug in -1 and then that would be the slope and then put the slope into either point slope or slope intercept form to get the equation of the tangent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That should give you your answer

OpenStudy (anonymous):

great!!! thank you so much.. one more question if you could help me out and do not mind.. how do i find the d/dx of y= 2(sqrt(x)) - (1/x)

OpenStudy (anonymous):

\[y=2\sqrt{x}-1/x\] This?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

oh actually its\[y = 2\sqrt{x} - 1/\sqrt{x}\]

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

The easiest way I find is to convert the sqrt into fractions.

OpenStudy (anonymous):

ok once i have it in a fraction how do i proceed?

OpenStudy (anonymous):

You differentiate the sum term by erm and factor out the constants

OpenStudy (anonymous):

So 2(d/dx(sqrt(x)))-d/dx (1/Sqrt(x))

OpenStudy (anonymous):

So you break them up into pieces.

OpenStudy (anonymous):

ok. so then the first term, would then end up as? 2(1/2 x^-1/2) which simplifies to x^-1/2

OpenStudy (anonymous):

well 1/Sqrt(x)= -1/2x^(3/2)

OpenStudy (anonymous):

yes your right on the first term

OpenStudy (anonymous):

so then would it simplify down to y1= x^(-1/2) - 1/2x^(3/2)

OpenStudy (anonymous):

The cleanest way to write it would be \[\frac{ 2x+1 }{ 2x^\frac{ 3 }{ 2 } }\]

OpenStudy (anonymous):

where did you get the 2x+1

OpenStudy (anonymous):

\[\frac{ 1 }{ 2x^ \frac{ 3 }{ 2 } }+ \frac{ 1 }{ \sqrt{x} }\]

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

np